OpenStudy (anonymous):

How to solve this ? If in any decreasing arithmetic progression , sum of all its terms,except the first term is equal to -36 and the sum of all its terms, except the last term is zero and the difference of the 10th and the 6th term is equal to -16, then first term of the series is ?

OpenStudy (kc_kennylau):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)

OpenStudy (kc_kennylau):

Let the first term be \(a\) and the common difference be \(d\) and the number of terms be \(n\) :)

OpenStudy (anonymous):

ok...

OpenStudy (kc_kennylau):

So the series will be: \(a\), \(a+d\), \(a+2d\), ..., \(a+(n-1)d\)

OpenStudy (kc_kennylau):

And the sum will be \(na+\dfrac{(n-1)n}2d\)

OpenStudy (kc_kennylau):

Now you can set 3 equations :)

OpenStudy (anonymous):

just a min plz ......

OpenStudy (kc_kennylau):

ok :)

OpenStudy (anonymous):

so which 2 equations out of 3 should i solve ?

OpenStudy (kc_kennylau):

what are your 3 equations?

OpenStudy (anonymous):

(n-1)d/2 = -36 , na = 0 , a+9d - ( a+5d ) = -16

OpenStudy (anonymous):

so , is a 0 ?

OpenStudy (kc_kennylau):

well your equations are not the same as mine...

OpenStudy (anonymous):

i am wrong anywhere ?

OpenStudy (anonymous):

sorry for disturbing u ! ...

OpenStudy (kc_kennylau):

Here are my equations: \[\left\{\begin{array}{} na+\frac{(n-1)n}2d-a &= &-36 &\mbox{(The first term is }a) \\ na+\frac{(n-1)n}2d-[a+(n-1)d] &= &0 &\mbox{(similar to above)}\\ (a+9d)-(a+5d) &= &-16 \end{array}\right.\]

OpenStudy (anonymous):

thanks a lot !

OpenStudy (anonymous):

i will solve it and further get back if I have any difficulty !

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

sorry once again for disturbing u !

OpenStudy (kc_kennylau):

it's not disturbing me at all :)

OpenStudy (anonymous):

gr8 of u !

OpenStudy (kc_kennylau):

done? :)

OpenStudy (anonymous):

i have got d = - 4 !

OpenStudy (kc_kennylau):

yep, then you can find n by subtracting (2) to (1) :)

OpenStudy (anonymous):

n=8 !

OpenStudy (anonymous):

i have got it now !

OpenStudy (anonymous):

Thanks a lot ! I can solve it further !

OpenStudy (kc_kennylau):

no problem :D Glad to know that I've helped you :)

OpenStudy (anonymous):

i have a question more !

OpenStudy (anonymous):

can i post it here ? or in a new post ?

OpenStudy (kc_kennylau):

please open a new post :)

OpenStudy (anonymous):

sure